WARNING: This web is not used anymore. Please use PDBService.MinuteS07November07 instead!
 

Minutes 7 November 07

  • Phone:
    • NDGF: Olli
    • PIC: Luis
    • TRIUMF: Andrew
    • SARA: Alexander
    • BNL: Carlos and John
    • GRIDKA: Andreas and Doris
    • CNAF: Barbara

  • CERN:
    • 3D: Dirk, Eva
    • ATLAS: Gancho, Florbela
    • LHCb: Marco

  • General announcements:
  • Workshop is coming, some changes ongoing. Please check the agenda.
  • Licenses predictions from several sites. Please keep track of the number of licenses at each site.
  • Dirk will send a spreadsheet to maintain the number of licenses (like last year).
  • Maria and Dirk will be next week at the OOW.
    • Co-joint presentation about Streams with Patricia McElroy, Streams product manager
    • Discussion with Oracle the still open Service Requests

  • 11g tests using Streams
    • Some promising results with new features but still some bugs to be fixed

  • Security patching: all CERN databases have been patched without problems

  • IT reorganization: 1st January we will be moved to new group called Data Management together with Castor and Grid development. Alberto Pace will be the group leader.

  • Sites status:
    • NDGF (Olli): storage intervention for next Tuesday. No schedule for the Cluster at NDGF
    • SARA (Alexander): some problems observed with ASM. Network intervention last week, expected to be transparent from outside.
    • GridKA (Doris): FTS setup finished, CPU patch to be applied – end November, beginning December
      • Doris will send around a draft UK presentation
    • Triumf (Andrew): CPU patch will be applied next week
      • Suggested to open a wiki page to report problems or send an email when there is a problem so sites can know what is happening.
    • PIC (Luis): new hw for next week. CPU patch scheduled for next week.
    • BNL (Carlos): CPU patch has been applied. One node was reconfigured (used by TAGs tests) in order to be used by the Streams failover during the patch apply.
      • New ASM diskgroup configured to host the TAG data, excluded from backups.
      • Proposal for the workshop: How to migrate from 32 to 64 bits? - Carlos volunteers for preparing this topic
    • CNAF (Barbara): new hw received, dual core servers, ATLAS cluster will be moved
      • Last week, problem with firewall settings. Now it is fixed.
      • CPU patch will applied next week.
      • VOMS tests run locally and successfully. No latency observed. From developers point of view everything went ok.

  • Experiment status:
    • ATLAS:
      • Streams replication fine
      • PVSS tests: PVSS schema moved to ATONR, filter rules to be defined because there are tables with user-defined types which are not supported by Streams. Average around thousand LCRs
      • TAGs: stress tests BNL and Triumf presented at the ATLAS week. Results in a wiki page, Florbela will send it around. Because hw configuration: differences on cpus number, memory, ... BNL showed double throughput than CERN integration and Triumf throughput is less than CERN integration
      • No changes on the Conditions data volume
      • Some changes on the TAGs data volume. Florbela will circulate the numbers.
    • LHCb:
      • Next week tests to limit number of connections
      • Jean-Philippe: sw changes done, now in sw certification, it will take at least 2 weeks more.
    • CMS (Dave and Eric):
      • Pool caching – discussion with Giacomo
Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2007-11-14 - EvaDafonte
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    PSSGroup All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback